# Bhubaneswar, India

### v1

@@ Abstract
Keywords:
Multi-Agent System Large language Models Simulation Orchestration Modularity
@@ 1 Introduction
The growing ubiquity of intelligent agent systems in domains such as autonomous robotics, collaborative software services, and multi-modal AI platforms has brought renewed focus on the challenge of effective task allocation [2]. In these systems, a central orchestrator is often responsible for selecting the most appropriate agent to handle a given task. The complexity of this decision stems from the heterogeneity of agents in terms of skills, contextual expertise, reliability, and adaptability — as well as from the inherent ambiguity and variety in real-world tasks [1]. Traditional approaches to agent orchestration often rely on static heuristics, fixed rules, or random selection strategies [12] . However, these methods struggle in dynamic environments where task requirements and agent capabilities evolve over time. Furthermore, many orchestration pipelines lack a robust mechanism to evaluate the quality of agent responses in a domain-agnostic and interpretable way, which hinders their ability to adapt and improve [3] .
@@ 2 System Architecture
MetaOrch is designed as a modular and extensible orchestration framework that facilitates intelligent agent selection in dynamic multi-agent environments. The architecture comprises five core components: The architecture comprises five core components: (1) Task Generator and Representation Module, (2) Agent Profile and History Tracker, (3) Neural Orchestration Model, (4) Fuzzy Evaluation Module, and (5) Supervised Learning Feedback Loop. Each component contributes to the overall pipeline, from interpreting incoming tasks to selecting the most appropriate agent, and then learning from the outcome to improve future decisions.
@@ 2.4 Fuzzy Evaluation Module
These are then combined using fixed, user-defined weights (e.g., completeness: 0.4, relevance: 0.4, confidence: 0.2) to yield a final fuzzy quality score.
These scores serve two purposes: (1) providing runtime feedback to human overseers, and (2) generating soft supervision signals to update the orchestration model in a self-supervised fashion [13] .
@@ 3.1 Agent Architecture
An agent’s performance on a task $\mathbf{t}$ with context $\mathbf{c}$ is deterministically cached and computed as:
@@
where $\epsilon_{i}\sim\mathcal{N}(0,1-r_{i})$ is the task-specific noise, and the final score is transformed into fuzzy evaluation metrics (completeness, relevance, confidence) [14].
@@ 4.2 Hyperparameter Optimization
The best configuration achieved 91.1% selection accuracy with a two-layer network, no dropout, learning rate of 0.01, batch size of 128, and confidence weight of 0.2.
@@ 4.3 Evaluation Results
MetaOrch achieved 86.3% selection accuracy, significantly outperforming all baselines. While Static-Best achieved high average quality (0.751), its low selection accuracy (5.7%) indicates it lacks contextual awareness for optimal agent-task matching.

### v2

@@ 1 Introduction
The growing ubiquity of intelligent agent systems in domains such as autonomous robotics, collaborative software services, and multi-modal AI platforms has brought renewed focus on the challenge of effective task allocation [10, 1]. In these systems, a central orchestrator is often responsible for selecting the most appropriate agent to handle a given task. The complexity of this decision stems from the heterogeneity of agents in terms of skills, contextual expertise, reliability, and adaptability — as well as from the inherent ambiguity and variety in real-world tasks [6]. Traditional approaches to agent orchestration often rely on static heuristics, fixed rules, or random selection strategies [15] . However, these methods struggle in dynamic environments where task requirements and agent capabilities evolve over time. Furthermore, many orchestration pipelines lack a robust mechanism to evaluate the quality of agent responses in a domain-agnostic and interpretable way, which hinders their ability to adapt and improve [2] .
@@ 2 System Architecture
MetaOrch is designed as a modular and extensible orchestration framework that facilitates intelligent agent selection in dynamic multi-agent environments.The architecture comprises five core components: (1) Task Generator and Representation Module, (2) Agent Profile and History Tracker, (3) Neural Orchestration Model, (4) Fuzzy Evaluation Module, and (5) Supervised Learning Feedback Loop. Each component contributes to the overall pipeline, from interpreting incoming tasks to selecting the most appropriate agent, and then learning from the outcome to improve future decisions.
@@ 2.4 Fuzzy Evaluation Module
These are then combined using fixed, user-defined weights (e.g., completeness: 0.4, relevance: 0.4, confidence: 0.2) to yield a final fuzzy quality score. In Eqs. (1–3), the variable score refers to the agent’s performance score as defined in Eq. (4) of Section 3.1. The confidence regression loss is computed as mean squared error between predicted and observed confidence scores. We performed a sensitivity analysis by varying the weights (0.4/0.4/0.2) and observed that selection accuracy fluctuated by ±2These scores serve two purposes: (1) providing runtime feedback to human overseers, and (2) generating soft supervision signals to update the orchestration model in a self-supervised fashion [8] .
@@ 3.1 Agent Architecture
Let d denote the dimensionality of the skill and task vectors, and c the dimensionality of the context vectors. An agent’s performance on a task $\mathbf{t}$ with context $\mathbf{c}$ is deterministically cached and computed as:
@@
where $\epsilon_{i}\sim\mathcal{N}(0,1-r_{i})$ is the task-specific noise, and the final score is transformed into fuzzy evaluation metrics (completeness, relevance, confidence) [16]. All vectors are sampled from a standard normal distribution with domain effects injected by adding a fixed bias to selected dimensions. For example, emergency tasks boost the first two skill dimensions by +1.0.
@@ 4.2 Hyperparameter Optimization
The best accuracy of 91.1% was achieved on the validation set during hyperparameter tuning, while the main result of 86.3% reflects test set performance on 300 held-out tasks. We acknowledge that 300 tasks is a limited sample and plan to expand to larger test sets, report results over multiple random seeds, and include confidence intervals and significance tests in future work
@@ 4.3 Evaluation Results
MetaOrch achieved 86.3% selection accuracy, significantly outperforming all baselines. While Static-Best achieved high average quality (0.751), its low selection accuracy (5.7%) indicates it lacks contextual awareness for optimal agent-task matching. To strengthen our evaluation, we plan to include additional baselines such as contextual bandits (e.g., LinUCB, Thompson Sampling), learned selectors (logistic regression, MLP), and multi-agent RL policies. This will provide a more comprehensive comparison against state-of-the-art orchestrators